Solve each of the following by substitution
5x-2y=-5
y-5x=3
Thanks for any help!!!
Heather

Answer by vidhyak(98)

5x-2y=-5
y-5x=3

Rearranging the equation
5x -2y = -5 .......eq 1
-5x +y = 3 ....... eq 2

Adding Eq1 and eq2 you get
-y = -2
y = 2

Substitute y = 2 in eq 1 oe eq 2
5x -2y = -5
5x -2(2) = -5
5x -4 = -5
5x = -1
x = -1/5